OpenStudy (anonymous):

If a radioisotope has a half life of 10 years, how much of the original material would remain after 40 years? A. One-eighth B. One-tenth C. One-half D. One-sixteenth

ganeshie8 (ganeshie8):

half life of 10 years, means, the quantity becomes half in 10 years

ganeshie8 (ganeshie8):

After 10 years ---- > \(\large \frac{1}{2} \)

OpenStudy (anonymous):

So answer C?

ganeshie8 (ganeshie8):

After 20 years ---- > \(\large \frac{1}{2} \times \frac{1}{2} \)

ganeshie8 (ganeshie8):

After 30 years ---- > \(\large \frac{1}{2} \times \frac{1}{2} \times \frac{1}{2} \)

ganeshie8 (ganeshie8):

After 40 years ---- > ?

OpenStudy (anonymous):

Oh i see. thank you

ganeshie8 (ganeshie8):

np :) every 10 years, it becomes half of watever it was before... it goes on reducing like that